### v1

# CER-HV: A CER-Based Human-in-the-Loop Framework for Cleaning Datasets Applied to Arabic-Script HTR

Abstract
Handwritten text recognition (HTR) for Arabic-script languages still lags behind Latin-script HTR, despite recent advances in model architectures, datasets, and benchmarks. We show that data quality is a significant limiting factor in many published datasets and propose CER-HV (CER-based Ranking with Human Verification) as a framework to detect and clean label errors. CER-HV combines a CER-based noise detector, built on a carefully configured Convolutional Recurrent Neural Network (CRNN) with early stopping to avoid overfitting noisy samples, and a human-in-the-loop (HITL) step that verifies high-ranking samples. The framework reveals that several existing datasets contain previously underreported problems, including transcription, segmentation, orientation, and non-text content errors. These have been identified with up to 90% precision in the Muharaf and 80–86% in the PHTI datasets.
We also show that our CRNN achieves state-of-the-art performance across five of the six evaluated datasets, reaching 8.45% Character Error Rate (CER) on KHATT (Arabic), 8.26% on PHTI (Pashto), 10.66% on Ajami, and 10.11% on Muharaf (Arabic), all without any data cleaning. We establish a new baseline of 11.3% CER on the PHTD (Persian) dataset. Applying CER-HV improves the evaluation CER by 0.3–0.6% on the cleaner datasets and 1–1.8% on the noisier ones. Although our experiments focus on documents written in an Arabic-script language, including Arabic, Persian, Urdu, Ajami, and Pashto, the framework is general and can be applied to other text recognition datasets.
keywords:
handwritten text recognition, data cleaning, Arabic scripts, label noise detection, CRNN, Pashto, Urdu, Persian, Ajami
journal: Pattern Recognition Journal
[…]
[label1] organization=Luleå University of Technology, Department of Computer Science, Electrical and Space Engineering, Embedded Intelligent Systems LAB, city=Luleå, postcode=97187, country=Sweden
1 Introduction
Handwritten text documents are a key part of cultural heritage and preserve historical knowledge across generations. Handwritten Text Recognition (HTR) aims to automatically convert images of handwritten documents into machine-readable text. This supports the preservation, search, and analysis of valuable materials and enables applications such as digitizing of archives or assisting individuals with visual impairments [11].
This work focuses on the recognition of handwritten text at the line level, where the goal is to predict the character sequence contained in a single extracted line of text. Recent work highlights this level of annotation as a practical and widely adopted choice in HTR, offering a balance between annotation effort and the ability to capture meaningful contextual information [18]. Compared to character- or word-level tasks, line-level recognition is more challenging and better reflects real document conditions [11].
HTR for Latin-script languages has reached near-human performance, especially for English [11]. In contrast, the same level of success has not been achieved for Arabic-script languages, which are used by more than 400 million people worldwide [30]. This gap is often attributed to the characteristics of Arabic script, which is cursive in nature, many letters differ only by small diacritical marks, and their shapes change depending on their position within a word, and available datasets are fewer, smaller, and less diverse than Latin-script resources [31]. To reduce this performance gap, research has mostly focused on architectural improvements while paying much less attention to data quality [6, 2].
In computer vision, the impact of noisy labels has been studied extensively, especially for classification tasks [35], but it has received almost no attention in text recognition. Our study shows that several widely used Arabic-script HTR datasets contain previously unreported issues that directly affect both training and evaluation.
These issues do not arise randomly; they often originate from the semi-automated way in which handwritten text datasets are constructed, where automated methods are combined with human oversight. Because producing large datasets is time and labor intensive, this oversight is not always adequately managed, allowing samples with errors—such as incorrect or incomplete transcriptions, segmentation problems (e.g., truncated lines or multiple text lines within a single image), orientation mistakes, script-mismatched samples, or non-textual document elements (e.g., stamps or signatures)—to be included in the final dataset. Label errors are problematic as deep neural networks can memorize noisy labels during training [41], which harms generalization [35] and can distort model selection and benchmarking [24].
Prior work on learning-dynamics-based noise detection shows that neural networks tend to learn clean samples first, then memorize harder or noisy ones [3, 34]. Methods such as O2U-Net [15] leverage this by ranking samples based on training loss, but these approaches were designed for classification tasks that use cross-entropy loss. In Connectionist Temporal Classification (CTC) based HTR, the loss is computed by summing the probabilities of all valid alignments (marginalizing over all paths) [12]. As a result, per-sample loss is influenced by alignment uncertainty and sequence length, making it less reliable for identifying mislabeled samples. Consequently, loss-based ranking is less informative for noise detection in HTR, motivating our use of Character Error Rate (CER) instead.
Building on this insight, we adapt learning-dynamics-based noise detection to HTR through three key design choices: replacing loss-based ranking with CER, removing cyclic retraining, and introducing early stopping to mitigate the effects of alignment uncertainty on learning dynamics in CTC-based sequence recognition, this yields an effective and practical noise detector for text recognition tasks such as HTR.
However, automated scoring alone can mistakenly flag valid but difficult samples as noisy, as prior work has shown that only about half of algorithmically flagged candidates correspond to genuine label errors [24]. Given the relatively small size of typical HTR datasets (e.g., IAM $\approx 10K$ lines [22], KHATT $\approx 6K$ [20]), human inspection of high-CER samples is practical. We therefore introduce CER-based Ranking with Human Verification (CER-HV), as a two-stage framework that combines CER-based ranking with human verification. To our knowledge, no prior work has combined noise detection with human verification in sequence-to-sequence (Seq2Seq) settings such as HTR.
Using this framework, we identified label errors across the training, validation, and test splits of five Arabic-script datasets. Although the overall proportion of errors is small in NUST-UHWR [36], KHATT [20], and PHTI [16], higher concentrations are observed in the Muharaf [30] and Ajami [40] datasets, where mislabeled samples constitute a non-negligible portion of the validation and test splits. Figure 1 presents examples of these error types.
Figure 1: Representative examples of error categories found in the datasets.
Overall, this study shows that data quality plays a major role in Arabic-script HTR and should be considered alongside model design. The proposed CER-HV framework helps identify and address label errors, improving the reliability of benchmark datasets in HTR. The main contributions of this work are summarized as follows:
  1. 1.
    We present the first systematic analysis of label and content errors in Arabic-script HTR datasets, defining a practical error taxonomy spanning transcription, segmentation, orientation, script mismatch, and non-text content.
  2. 2.
    We introduce CER-HV, a CER-based ranking with human verification framework for HTR that adapts learning-dynamics-based noise detection to CTC-based sequence recognition through CER-based scoring and early stopping. This can be used as a general procedure for inspecting and validating label quality in line-level HTR datasets.
  3. 3.
    We demonstrate that data quality significantly distorts reported HTR benchmarks, and quantify the impact of label noise across six Arabic-script datasets, showing evaluation CER reductions of up to 1.8% and establishing updated reference results on cleaned evaluation splits.
  4. 4.
    We establish a strong CRNN baseline for Arabic-script HTR that achieves state-of-the-art results on multiple datasets, including a reduction from 20.7% to 8.26% CER on Pashto (PHTI), and the first benchmarks for Ajami and Persian line-level HTR.
  5. 5.
    We provide the first publicly available cleaned evaluation splits and line-level benchmarks for Persian handwritten text (PHTD), including extracted line images and test partitions, enabling reproducible comparison and future benchmarking.
  6. 6.
    All code, cleaned splits, and human-verified error annotations are released to support reproducible and reliable benchmarking in Arabic-script HTR [^1] https://github.com/SanaNGU/CER-HV.
The remainder of this paper is organized as follows. Section 2 reviews related work on HTR for Arabic-script languages and prior research on learning with noisy labels. Section 3 describes the proposed two stage noise detection framework and explains how CER and learning dynamics are adapted to sequence prediction. Section 4 presents the experimental setup, including the datasets, preprocessing, and training details. Section 5 reports the evaluation results and the effect of cleaning label errors on CER. Section 6 concludes the paper.
2 Related Work
2.1 Arabic-script HTR
Arabic-script HTR remains challenging because the script is inherently cursive and context-dependent: the same letter may take different shapes depending on whether it appears in isolated, initial, medial, or final position, and many letters differ only by dots or optional diacritics (tashkīl/ḥarakāt), which are small and easily missed in degraded scans. These properties complicate segmentation and increase confusion among visually similar characters, especially in historical material with diverse calligraphic styles and layout noise [30, 31, 40].
Early Arabic-script HTR work utilized hidden markov models and handcrafted features, focusing primarily on isolated word or subword recognition within closed-set lexicons, such as the IFN/ENIT dataset—a benchmark consisting of Tunisian town and city names [9, 10, 4]. While these systems achieved high precision on fixed vocabularies, they relied on explicit or semi-explicit segmentation assumptions, such as predefined sliding windows, grapheme hypotheses, or lexicon-driven alignment. This reliance limited their robustness to unconstrained handwriting and hindered effective generalization to open-vocabulary, line-level transcription.
With the shift to deep learning, Arabic-script HTR increasingly moved toward segmentation-free sequence modeling that maps visual feature sequences directly to text, combining convolutional neural networks (CNNs) for feature extraction with recurrent sequence models (e.g., BLSTM, GRU, or MDLSTM) and training them using CTC [12]. A key motivation for adopting CTC is that, in HTR, the alignment between the input image sequence and the output character sequence is generally unknown; CTC enables end-to-end training by marginalizing over all valid alignments without requiring explicit frame-level or character-level segmentation. Within this paradigm, CRNN-style architectures based on a CNN–RNN–CTC design, as introduced by Shi et al. [32], became a common baseline, as they jointly learn visual representations and sequential dependencies while avoiding explicit segmentation. Variants of this design, often strengthened through deeper convolutional backbones, data augmentation, or architectural refinements, were subsequently adopted across multiple Arabic-script languages [30, 36, 5].
In line with developments in Latin-script HTR, Arabic-script research has increasingly explored Transformer-based architectures to replace or augment recurrence. By leveraging self-attention [37], these models offer alternatives to CTC through sequence-to-sequence (Seq2Seq) or transducer-style decoding [23]. For example, HATFormer adapted the English-centric TrOCR model to historical Arabic manuscripts, demonstrating that architectures built for Latin scripts can be transferred to Arabic with appropriate preprocessing and training strategies [6].
A similar direction is observed in Urdu HTR, where Transformer-based architectures have also been explored [28], often relying on additional printed or synthetic text data to supplement handwritten training sets due to their data-intensive nature [21, 14, 13].
Another Arabic-script language that has received comparatively limited attention at the line level is Persian. Much of the existing work on Persian HTR focuses on isolated characters, digits, or word-level classification [29], largely due to the absence of publicly available line- or page-level Persian handwriting datasets. This motivates the use of the PHTD dataset [1] and the extraction of line images to enable systematic benchmarking of Persian HTR at the line level.
Pashto represents another low-resource case within Arabic-script languages, where progress at the line level has long been constrained by limited data availability. Until recently, the lack of suitable datasets restricted most research to isolated characters or small-scale experiments. The introduction of the Pashto Handwritten Text Imagebase (PHTI) [16] enabled the first large-scale benchmark on real-world Pashto handwritten text lines, where performance was evaluated using an MD-LSTM-based model [17]. However, evaluation has largely remained confined to this single dataset, with limited comparative analysis across different recognition models.
Ajami HTR is an even more recent research direction. The first publicly available Ajami HTR dataset was introduced only recently [40], providing manually segmented and transcribed manuscripts for West African languages written in adapted variants of the Arabic script, specifically Hausa and Fulfulde. While this work demonstrates that existing Arabic-script recognition models perform poorly on Ajami manuscripts, it does not establish standardized benchmarks or comparative baselines. As a result, Ajami remains largely unexamined in systematic HTR evaluations, further motivating its inclusion in multi-language benchmarking studies.
A key gap across this literature is that Arabic-script HTR is still commonly pursued per language/script community in isolation (Arabic vs. Persian vs. Urdu vs. Ajami, etc.), despite substantial overlap across these languages. As a result, the extent to which a single, well-configured recognition model exhibits consistent behavior across different Arabic-script languages has not been systematically examined. We also show that such cross-language behavior can be assessed by benchmarking the same CRNN architecture across multiple Arabic-script datasets.
2.2 Label Error Detection in Handwritten Text Recognition
Learning with noisy labels has been extensively studied in machine learning, primarily in the context of image classification. Comprehensive surveys show that most existing approaches are designed for fixed-label prediction problems trained with cross-entropy loss, and can be broadly grouped into loss-robust training, sample selection, label correction, and noise-transition modeling strategies [35]. Within this literature, label error detection is typically achieved by exploiting training dynamics (e.g., small-loss behavior), prediction confidence, or agreement across models.
In contrast, label error detection has received comparatively little attention in HTR and related Seq2Seq text recognition tasks. Unlike classification, HTR models predict variable-length character sequences and are trained using sequence-level objectives, including alignment-based losses such as CTC as well as cross-entropy losses in encoder–decoder Transformer architectures [11]. As a result, the training signal reflects alignment uncertainty and decoding effects in addition to transcription correctness, which makes the direct adoption of classification-oriented noise detection strategies less straightforward and limits the extent to which conclusions from the classification literature can be transferred to HTR without adaptation.
Several prominent noise detection approaches rely on model confidence estimates [35]. Confident learning, for example, identifies potential label errors by comparing predicted class probabilities against the given labels under a class-conditional noise assumption [25]. While such methods are effective for classification, their extension to sequence prediction tasks presents additional challenges, as confidence must be defined over entire sequences rather than individual classes and is often miscalibrated in practice. In particular, studies in Optical Character Recognition (OCR) report that word-level confidence scores do not consistently correlate with character or word error rates and depend strongly on OCR system settings [8]. Similarly, recent work on scene-text recognition demonstrates that modern sequence models are systematically overconfident, and that raw confidence scores do not reliably reflect true transcription accuracy [33]. These findings indicate that confidence-based noise detection, while well-founded in classification, remains challenging to apply reliably in Seq2Seq text recognition tasks such as HTR.
Another family of methods exploits learning dynamics to identify noisy labels. O2U-Net ranks samples based on their average training loss over cycles of underfitting and overfitting, under the observation that mislabeled samples tend to maintain higher loss throughout training [15]. This idea is closely related to the memorization effect, where deep networks fit clean samples earlier than noisy ones. This early-learning phenomenon has been analyzed more broadly in the context of learning with noisy labels, where several works show that deep networks first fit clean labels before memorizing noisy ones, and propose training strategies that explicitly exploit or regularize this behavior to improve robustness [19, 38]. While O2U-Net is formulated for classification, its reliance on per-sample loss trajectories makes it conceptually applicable to other settings. However, in HTR, the interpretation of per-sample CTC loss is complicated by alignment variability and structural errors, and cyclic overfitting strategies may not always be desirable. Prior work on early stopping shows that preventing late-stage memorization can be beneficial under label noise, suggesting that careful control of training dynamics is important when adapting such methods [34].
Overall, the literature suggests that while label error detection is well studied for classification, there is limited work that directly addresses noisy supervision in Seq2Seq text recognition, including HTR, OCR, and scene-text recognition. This gap motivates strategies for identifying samples with errors in HTR, forming the basis of the approach explored in this work.
3 Method
In this section, the proposed cleaning approach is described. As illustrated in Figure 2, the framework is divided into two main stages. The first stage is an automatic procedure that scores each sample based on the CER produced by a CRNN using the initial sample labels. The second stage is a human-in-the-loop (HITL) procedure that verifies the highest ranking samples. The goal is to identify mislabeled or out-of-scope samples and improve the quality of both the labels and the image content in HTR datasets. The two stages together form the CER-HV framework, which stands for CER-based Ranking with Human Verification. The following subsections present the problem formulation, the CRNN model used for recognition, and the details of the CER-HV framework.
Figure 2: Overview of the CER-HV framework. The first stage employs the learning dynamics of the CRNN and scores each sample using CER, and the second stage applies a human verification step to the highest ranking samples ($CER>\tau$).
3.1 Problem Formulation
The task of HTR is to map an image of handwritten text to a sequence of characters. Let $\mathcal{X}$ denote the space of text line images in the current dataset, where each $x\in\mathcal{X}$ is a tensor in $\mathbb{R}^{H\times W\times C}$ with height $H$, width $W$, and channels $C$. Let $\Sigma$ be the alphabet of characters, and let $\Sigma^{\ast}$ be the set of all possible sequences over this alphabet.
For an input $x$, with ground-truth transcription $y\in\Sigma^{\ast}$, the goal of handwritten text recognition is to predict the most likely character sequence:
$y^{\ast}=\arg\max_{y\in\Sigma^{\ast}}P(y\mid x)\,.$ (1)
In CTC based models such as the CRNN, the network outputs a sequence of frame level probability distributions over the extended alphabet $\Sigma^{\prime}=\Sigma\cup\{\text{blank}\}$. The blank symbol allows the model to handle variable length alignments by representing either no character or repeated characters. Let $T$ be the length of the output sequence produced by the encoder. Each alignment path $\pi=(\pi_{1},\dots,\pi_{T})$ belongs to $(\Sigma^{\prime})^{T}$ and collapses to a final transcription through a function $B$ that removes blanks and merges repeated characters.
The probability of $y$ is defined as the sum over all valid alignment paths:
$P(y\mid x;\theta)=\sum_{\pi\in(\Sigma^{\prime})^{T},B(\pi)=y}\prod_{t=1}^{T}P(\pi_{t}\mid x;\theta)$ (2)
and the model parameters $\theta$ are learned by minimizing the CTC loss over the training set
$\theta^{\ast}=\arg\min_{\theta}\mathcal{L}_{\mathrm{CTC}}(\theta;D)\,.$ (3)
The dataset $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ may contain noisy labels. Let $y_{i}^{\ast}$ denote the unknown true label and $z_{i}\in\{0,1\}$ indicate whether the observed label is noisy. Noise in text recognition can come from transcription mistakes, segmentation problems, orientation issues, or irrelevant and out-of-scope content.
The CER-HV framework aims to identify samples that are likely to have label noise. Given the model prediction, obtained as
$\hat{y}_{i}=\text{Decode}(f_{\theta}(x_{i}))$ (4)
where $f_{\theta}$ denotes the recognition model, each sample is assigned a score
$c_{i}=\mathrm{CER}(\hat{y}_{i},y_{i})$ (5)
where CER computes the normalized edit distance between the predicted and observed sequences. Although high CER values often indicate potential label errors, they can also arise from correctly labeled but visually difficult samples. For this reason, samples with high CER values are ranked as likely to have $z_{i}=1$ and are later verified in the human verification stage.
3.2 HTR Model: CRNN Architecture
We instantiate the recognition model $f_{\theta}$ in Eq. 4 using a CRNN for line-level HTR. While any well-performing deep learning recognizer could be used to identify potential label errors, we employ a CRNN which, based on our experience, performs robustly across all Arabic-script datasets evaluated in this study. The CRNN was selected for two additional reasons: its computational efficiency with fewer parameters than Transformer-based models, and the opportunity to establish the Best Practices CRNN [27] as a strong baseline for Arabic-script HTR.
The CRNN model performs handwritten text line recognition using CTC [12], eliminating the need for explicit character-level segmentation. Given an input image of size H×W (height × width), the network interprets it as a sequence of W narrow vertical slices, each representing a small visual context of the image. The model outputs a sequence of character probability distributions of length L, typically shorter than W due to spatial downsampling. During decoding, consecutive duplicate predictions are merged and blank tokens are removed to yield the final transcription.
Our architecture builds upon the original CRNN framework proposed by Shi et al. [32], which integrates convolutional, recurrent, and transcription components into a unified end-to-end model. We adopt the refined implementation introduced in the Best Practices framework [27], which enhances the CRNN design with deeper residual layers, normalization, and improved training strategies. As illustrated in Figure 3, the model follows an encoder–decoder structure, consisting of a CNN encoder for visual feature extraction and an RNN decoder for sequence modeling and transcription.
Figure 3: The CRNN architecture used as the base recognition model in Stage 1 of the CER-HV framework. The model combines a convolutional encoder for visual feature extraction with an RNN decoder and a CTC output layer for sequence prediction.
The CNN encoder extracts hierarchical visual features from the input text line image, producing a sequence of 2D feature maps. An intermediate column-wise max-pooling step reduces the feature map height to one, effectively converting the 2D representation into a 1D sequence of feature vectors along the image width. This operation preserves the horizontal spatial order while reducing sensitivity to absolute vertical position, preparing the sequence for the recurrent decoder.
The RNN decoder, composed of three stacked bidirectional LSTM layers (256 hidden units each), models contextual dependencies across time steps in both directions. Each recurrent output is passed through a fully connected classifier layer that maps it to a vector of dimension equal to the character set size (including the CTC blank token). The resulting per-frame character probability distributions are then decoded by the CTC layer to produce the final recognized text sequence without requiring explicit segmentation.
While most Arabic-script HTR systems adopt CRNN structures combining convolutional feature extraction, recurrent sequence modeling, and CTC decoding, implementation details significantly affect performance. Earlier approaches such as Shi et al. [32] and Puigcerver [26] employed shallow VGG-style convolutional backbones that limited fine-grained stroke feature capture. Following the Best Practices framework [27], our implementation incorporates three critical modifications: (1) a deeper residual CNN encoder with batch normalization for improved feature learning, (2) column-wise max pooling instead of concatenation to reduce feature dimensionality while maintaining vertical invariance, and (3) an auxiliary CTC shortcut branch during training to enhance recurrent layer convergence.
Concretely, the CNN encoder begins with a 7×7 convolution (32 channels, stride 2×2), followed by ResNet blocks (2 with 64 channels, 3 with 128, and 2 with 256), interleaved with batch normalization and dropout. Progressive 2×2 max-pooling between block groups downsamples the spatial resolution, producing feature maps of size (H/8) × (W/8) × 256. Column-wise max pooling then flattens this tensor into a sequence of (W/8) feature vectors (256 dimensions each).
The recurrent head and decoding stage use three stacked BiLSTM layers (256 hidden units each) to process the feature sequence. A fully connected classifier projects each BiLSTM output onto the character space, followed by softmax normalization to obtain per-frame probability distributions. During inference, greedy CTC decoding selects the most probable character at each time step, merges duplicates, and removes blanks to generate the final transcription.
To improve convergence and discriminative feature learning, the model employs an auxiliary CTC shortcut branch trained jointly with a weighted multi-task loss (weight = 0.1). This shortcut is discarded during inference, introducing no additional computational overhead at test time.
This CRNN architecture is used in the CER-HV framework as the base recognizer $f_{\theta}$. Its predictions $\hat{y}_{i}=\text{Decode}(f_{\theta}(x_{i}))$ are used to compute the CER scores $c_{i}$ that guide the ranking of samples in the noise detection stage.
3.3 CER-HV Framework: CER Based Ranking and HITL Verification
The second component of the methodology focuses on improving data quality through automated label error detection and human verification. The CRNN recognizer depends critically on the quality of the training data. Potential label noise such as incorrect transcriptions, image segmentation errors, or irrelevant samples can significantly degrade accuracy [24]. To address this, we introduce CER-HV, a two-stage framework that combines automated label noise detection based on CER scoring with human verification. Figure 2 shows the overall flow.
Stage 1: Label Noise Detection
We adapt the O2U-Net approach [15], originally proposed for classification, to the Seq2Seq nature of HTR. O2U-Net leverages the learning dynamics of deep networks, where clean samples are learned early and noisy samples are fitted later. However, two key adaptations are required to make this approach effective for HTR.
While inspired by prior learning-dynamics-based work such as O2U-Net, CER-HV represents a reformulation of learning-dynamics-based noise detection for CTC-based sequence recognition, differing in key respects. First, we replace the original loss-based scoring with the CER, the standard evaluation metric in HTR. In classification tasks, the training loss correlates well with label correctness, but in Seq2Seq problems, the CTC loss measures alignment rather than transcription accuracy. By contrast, CER provides a direct and interpretable measure of prediction quality: for example, a CER of 35% indicates that 35% of characters differ from the ground truth, making it immediately meaningful for human reviewers.
Second, rather than using the cyclical multi-round training procedure of O2U-Net to avoid overfitting, we employ an early-stopping criterion. Prior studies have shown that early stopping based on validation metrics effectively detects convergence and prevents overfitting [34]. We identify a single, objectively defined convergence epoch $t_{\mathrm{conv}}$ at which training and validation CER values converge and further improvement becomes negligible. This choice simplifies the procedure, removes the need for hyperparameter tuning, and aligns better with sequence-prediction tasks such as HTR.
Once the CRNN reaches the convergence epoch, each training sample is assigned a score
$c_{i}=\mathrm{CER}(f_{\theta_{t_{\mathrm{conv}}}}(x_{i}),y_{i})\,.$ (6)
Here, $f_{\theta_{t_{\mathrm{conv}}}}$ denotes the recognition model evaluated at the convergence epoch $t_{\mathrm{conv}}$, determined by early stopping. The score $c_{i}$ measures the difference between the model prediction and the given label. Samples are then ranked in descending order of $c_{i}$, with higher values indicating a greater likelihood of label noise.
Stage 2: Human Verification
Although automated scoring highlights suspicious samples, some correctly labeled but visually difficult examples may also receive high CER. For this reason, a HITL verification step is used, in which samples whose CER exceeds a threshold are selected for review:
$\mathcal{S}=\{i\mid c_{i}>\tau\}\qquad\tau=0.25\,.$
The threshold $\tau=0.25$ was selected based on consistent observations across the five datasets that exhibited label noise, considering their training, validation, and test splits. Across these splits, samples with CER below 0.25 were overwhelmingly correctly labeled, while values above this threshold reliably indicated potential label or content errors. This value therefore represents a stable and practical operating point rather than a sensitive hyperparameter. In practice, $\tau$ simply controls review volume and can be adjusted based on available verification budget.
Human reviewers inspect the selected samples and assign them to one of the following categories:
  • 1.
    Transcription error– the ground truth text does not match the handwritten content in the image.
  • 2.
    Segmentation error– multiple text lines are included within a single image, or the lines are truncated/incomplete.
  • 3.
    Orientation error– significantly rotated text lines that contradict the expected reading direction.
  • 4.
    Script mismatch– the content is written in a script or language different from the target alphabet (e.g., Latin or numeric text appearing in Arabic-script datasets).
  • 5.
    Irrelevant or non-text samples– stamps, signatures or other non-textual document elements.
  • 6.
    Valid but hard– the sample is correct, but the sample is visually challenging to correctly automatically transcribe.
Human reviewers relabel, fix, or remove noisy samples while keeping the valid but difficult examples. This step mitigates systematic biases in automated ranking, particularly its tendency to flag valid but challenging samples as noisy.
After human verification and cleaning, the cleaned dataset $D^{\prime}$ is used to retrain the model. The retrained recognizer is evaluated on the test set, and the same CER based scoring procedure can be used to identify remaining issues in the evaluation data. The process may be repeated if needed. Algorithm 1 summarizes the complete procedure. Beyond cleaning existing benchmarks, CER-HV can be applied as a general validation procedure to assess label quality in newly created or previously unverified line-level HTR datasets. Training details, hyperparameters, input preprocessing, and experimental settings are described in Section 4.
Notation: $D=\{(x_{i},y_{i})\}_{i=1}^{N}$ denotes a dataset split (training, validation, or test), where $x_{i}$ denotes a text line image and $y_{i}$ denotes its ground-truth transcription; $f_{\theta_{t_{\mathrm{conv}}}}$ is the recognizer evaluated at the convergence epoch; $\hat{y}_{i}=\mathrm{Decode}(f_{\theta_{t_{\mathrm{conv}}}}(x_{i}))$ is the predicted transcription; $c_{i}=\mathrm{CER}(\hat{y}_{i},y_{i})$ is the CER score; $\mathcal{S}$ is the set of selected high-CER samples; $D^{\prime}$ is the cleaned dataset; $M$ is the final trained model; $T_{\max}$ denotes the maximum number of training epochs; $\theta$ denotes the trainable parameters of the recognizer; $\theta_{t}$ denotes the parameters after epoch $t$; $\tau$ denotes the CER threshold used to select samples for human inspection.
0.8em
Input: Training set $D=\{(x_{i},y_{i})\}_{i=1}^{N}$
Output: Cleaned set $D^{\prime}$, trained model $M$
Stage 1 : Label Noise Detection;
Initialize $\theta$, learning rate $\eta$, batch size $b$, early stopping patience $P$;
for $t=1$ to $T_{\max}$ do
  Train the CRNN on mini batches from $D$ using the CTC loss;
  Update parameters $\theta_{t}=\theta_{t-1}-\eta\nabla_{\theta}\mathcal{L}_{\mathrm{CTC}}$;
  Compute validation CER(t)(t);
  if validation CER does not improve for $P$ epochs then
     Set convergence epoch $t_{\mathrm{conv}}=t$;
     break;
Compute per sample CER at convergence:;
for each $i$ in $D$ do
  $c_{i}=\mathrm{CER}(f_{\theta_{t_{\mathrm{conv}}}}(x_{i}),y_{i})$;
Stage 2: Human Verification;
Rank all samples by descending $c_{i}$;
Select samples $\mathcal{S}=\{i\mid c_{i}>\tau\}$ for human inspection;
for each $(x_{i},y_{i})\in\mathcal{S}$ do
  Human inspects sample and determines error type:
 if label is incorrect then
    Relabel as $(x_{i},\tilde{y}_{i})$;
  else if sample has segmentation, orientation, or irrelevant content then
    Fix or remove sample;
  else
    Keep sample as valid but difficult;
Construct the cleaned dataset $D^{\prime}$ ;
Retrain the CRNN on $D^{\prime}$ with early stopping to obtain final model $M$;
while further corrections may remain do
  Optionally repeat Stages 1 to 2 on the updated $D^{\prime}$;
Algorithm 1 CER-HV Framework for Label Error Detection
4 Experimental Setup
This section describes the datasets and training setup used to evaluate the CER-HV framework.
4.1 Datasets
This study employs six Arabic-script HTR datasets covering five languages segmented and labeled at the line level, covering both modern and historical handwriting. Together, these datasets provide broad coverage of writing styles, domains, and regional variations across Arabic, Persian, Urdu, Pashto, and West African Ajami scripts. Examples from the datasets are shown in Figure 4, and dataset statistics are summarized in Table 1. For datasets without public train, validation, and test splits (e.g., PHTD and Ajami), we applied an 80%, 10%, and 10% division. These datasets collectively enable assessment of the robustness of the proposed framework across diverse handwriting styles and linguistic varieties.
Figure 4: Examples of handwriting from six datasets. From top to bottom: KHATT (Arabic), Muharaf (Arabic), PHTI (Pashto), PHTD (Persian), NUST-UHWR (Urdu), Ajami Hausa, and Ajami Fulfulde.
Table 1: Dataset descriptions including train/val/test split sizes, followed by average line image dimensions and character length ranges.
Dataset Language Train/Val/Test Avg. Width–Height Max Char Len
KHATT Arabic 4837/938/967 2075–200 132
Muharaf (public) Arabic 22091/1069/1334 575–60 181
PHTI Pashto 25200/5400/5400 992–77 180
PHTD Persian 1787 1960–178 197
NUST-UHWR Urdu 8483/1061/1061 950–64 116
Ajami Hausa + Fulfulde 6132 1290–215 106
KHATT (Arabic): The KHATT dataset [20] is a widely used benchmark for modern Arabic handwriting recognition. It contains paragraph images written by 1,000 writers from diverse demographic backgrounds. Each writer contributed both repeated and unique texts under controlled imaging conditions (black ink on white background). In our experiments, we used only the unique-text subset, yielding 6,742 segmented line images.
Muharaf (Arabic): The Muharaf dataset [30] comprises historical Arabic manuscripts digitized and manually transcribed at the line level. It features considerable variation in calligraphy styles, ink intensity, and document aging effects, making it particularly challenging for modern HTR systems. Our experiments use the public release containing approximately 24K line images.
PHTI (Pashto): The Pashto Handwritten Text Imagebase (PHTI) is a dataset containing 36,082 handwritten text-line images representing real-world Pashto text, along with ground truth annotations in UTF-8 encoding. The dataset encompasses diverse genres including short stories, historical narratives, poetry, and religious content, collected from a variety of modern writers.
PHTD (Persian): The PHTD dataset [1] consists of 140 handwritten Persian pages produced by 47 writers across three categories of content: historical documents, school dictations, and general writing. The original dataset provides pixel-level ground-truth masks for each page but does not include predefined line images or bounding boxes or train–validation–test splits.
To generate consistent and clean line images, we extracted all text lines using a mask-based cropping procedure that leverages the pixel-level annotations provided with the dataset. For each line, a binary mask was generated, the minimal bounding box enclosing all labeled pixels was computed with a small padding of 5 px, and the cropped region was masked to retain only the target line while setting all other pixels to background. This procedure ensures accurate isolation of individual lines despite the substantial overlap between bounding boxes caused by ascenders and descenders in cursive Persian handwriting.
Because several pages contain highly similar or near-duplicate content, we performed a similarity analysis to prevent data leakage between training and evaluation splits. Duplicate detection revealed that 2 pages were exact duplicates, and among the 9,730 possible page pairs, 1,155 pairs (11.9%) exhibited more than 85% textual similarity. Only 37 pages showed no significant overlap with any other page. The validation and test sets were selected exclusively from these 37 non-overlapping pages to ensure that no evaluation sample is textually related to the training data.
The final split contains 110 pages for training (1,473 line images), 15 pages for validation (160 lines), and 15 pages for testing (155 lines), corresponding to an 80%,10%,10% division applied at the page level to ensure a leakage-free evaluation protocol for Persian handwritten text recognition. The cleaned line images and the final train, validation, and test splits will be released publicly [^2] PHTD cleaned dataset: https://huggingface.co/datasets/sana-ngu/PHTD.
NUST-UHWR (Urdu): The NUST-UHWR dataset [36] is the first published large-scale, unconstrained Urdu handwriting corpus. It includes 10,608 text lines written by nearly 1,000 contributors. Text samples span seven topical domains to ensure linguistic variety. The dataset is partitioned into 8,483 training, 1,061 validation, and 1,061 test lines.
Ajami (Hausa and Fulfulde): The Ajami dataset  [40] is the first publicly released benchmark for HTR of West African manuscripts, comprising 29 manually segmented and transcribed documents written in Fulfulde and Hausa. The term “Ajami” originates from Arabic and means “foreign”; it refers to African languages written using the Arabic script. The initial corpus contains 6,132 line images, showcasing a significant challenge for HTR systems due to its diversity of script styles, including Maghrebi, Hausa Naskh, and the distinct Hausa Thuluth calligraphy.
Critically, to prevent data leakage between training and evaluation splits, we performed a similarity check and ensured that no samples were shared between the validation and test sets. The baseline Arabic HTR models used for evaluation by the dataset creators were trained on Middle Eastern and South Asian manuscripts, and yielded high CERs ranging from 65–84% when applied to this corpus. We prepared and publicly released the extracted line images and data splits used in this work as an open-source resource to ensure reproducibility for future research.
4.2 Training Setup and Evaluation
Training setup
All experiments were conducted on NVIDIA A100-SXM4-40GB GPUs using a CRNN architecture configured for Arabic-script handwriting.
For each dataset, image preprocessing followed a standardized resizing and padding procedure to reduce shape distortion and ensure uniform input dimensions. Specifically, all line images were resized to the rounded average height and width (in whole pixels) of their respective training set, as reported in Table 1. If an image exceeded these target dimensions, it was proportionally rescaled; otherwise, it was centered and padded with median pixel values. This approach preserved the aspect ratio while providing consistent spatial context across samples, which we observed stabilized training and improved recognition performance on longer text lines. In addition to resizing, we applied fixed horizontal padding of 64 pixels to the left and right sides of each line image, following a preprocessing strategy reported in prior work [7], which demonstrated improved recognition performance in CRNN-based line-level HTR.
To further regularize training and enhance robustness to handwriting variability, we applied a consistent set of data augmentations across all datasets. These include affine transformations (small rotation and skew), grid distortion, elastic deformation, Gaussian noise addition, brightness and contrast adjustment, and morphological operations (erosion and dilation). Such augmentations have been shown to improve generalization in HTR tasks by encouraging invariance to writer style, line curvature, and ink intensity.
For each dataset, a separate CRNN model was trained for a maximum of 800 epochs with a batch size of 16 using the Adam optimizer and an initial learning rate of $5\times 10^{-4}$. A multi-step learning rate scheduler reduced the learning rate by a factor of 0.1 at 50% and 75% of the total epochs. Early stopping with patience $P=20$ epochs was applied based on validation CER to prevent overfitting. For each dataset, the model achieving the lowest CER on the validation set was selected for final evaluation on the test partition.
Evaluation Metric
Recognition performance was assessed using the Character Error Rate (CER), the standard metric in HTR. CER is computed as the normalized Levenshtein (edit) distance between the predicted and ground-truth transcriptions, reflecting the proportion of character insertions, deletions, and substitutions required to match the prediction to the reference. Lower CER indicates higher transcription accuracy and provides a fine-grained measure of model performance at the character level.
5 Results
This section presents the experimental results of our study, focusing on baseline CRNN performance and the impact of dataset label quality. We also assess the effectiveness of the proposed CER-HV framework.
We begin by comparing our CRNN results with prior work across all datasets using the original, unaltered data to establish quantitative CER baselines. We then apply the CER-HV framework to analyze label noise across the training, validation, and test splits. Representative qualitative examples of common labeling issues are presented to illustrate the identified error categories. Finally, we quantify the impact of correcting these errors by evaluating how cleaning the affected splits influences recognition performance.
5.1 Quantitative Benchmarking of the CRNN-Based HTR System
The proposed CRNN architecture (Section 3.2) was evaluated across the six Arabic-script datasets described in Section 4.1 on the original data without any data cleaning. Table 2 summarizes the performance of this CRNN in comparison with previously reported results. For our CRNN results, we report mean CER with standard deviation across three independent training runs.
While most CNN–BiLSTM–CTC models share a similar high-level structure, implementation details can strongly influence recognition performance. As described in Section 3.2, our CRNN follows the Best Practices framework [27], which contributes to the consistent performance gains observed across datasets. On the KHATT dataset, our model achieves a new SoTA CER of 8.45%, outperforming both traditional CRNN-based methods [5] and recent Transformer based systems such as TrOCR [6] and DAN [2], despite not using synthetic data.
For Muharaf, we outperform the baseline model from [30] by eight absolute percentage points, reducing CER from 18.1% to 10.11%. Furthermore, our model surpasses the Transformer-based TrOCR system from [6], which reports 11.7% CER on Muharaf using both the public and restricted portions of the dataset as well as large-scale synthetic data. Without access to the restricted split and without relying on any additional data, our CRNN achieves 10.11% CER on the public subset alone.
For PHTI, our CRNN achieves a substantial improvement over prior non-transformer models, reducing CER from 20.7% to 8.26%, representing one of the largest performance gains across all evaluated datasets. For PHTD, where publicly reported benchmarks are limited, our model establishes a new baseline of 11.3% CER. On NUST-UHWR, our model reaches 6.66% CER, improving upon earlier convolutional and recurrent baselines [36, 28, 13], although Transformer-based architectures [21, 13] still achieve superior performance due to attention mechanisms and synthetic pretraining. For Ajami, the model attains 10.66% CER, far outperforming previously reported results (64-84%), which were obtained through zero-shot evaluation using pretrained OPEN-ITI models [40].
In summary, the results demonstrate that a carefully optimized CRNN, trained without transformer attention or synthetic augmentation, can match or surpass more complex architectures on multiple Arabic-script datasets.
Table 2: Comparison of our CRNN with established methods from the literature. Datasets, models, architectures, synthetic data usage, and CER results across datasets (reported as mean ± standard deviation). A dashed line separates non-Transformer architectures from Transformer-based models.
Dataset & Language Model Architecture Synthetic Data CER (%)
KHATT (Arabic)[20] [30] SFR Yes 14.1
[5] CNN+BiLSTM+CTC No 15.45
[5] EfficientNetB4BLSTM Yes 13.34
[39] FCN No 8.77
Ours CRNN No 8.45 ± 0.14
[23] Transformer with cross-attention Yes 18.45
[6] TrOCR Yes 15.4
[2] DAN No 8.9
Muharaf (Arabic) [30] [30] SFR Yes 18.1
Ours CRNN No 10.11 ± 0.055
[6] TrOCR Yes+11K restricted data 11.7
PHTI (Pashto) [16] [17] MD-LSTM No 20.7
Ours CRNN No 8.26 ± 0.05
PHTD (Persian) [1] Ours CRNN No 11.3 ± 0.07
NUST-UHWR (Urdu) [36] [36] Modified CRNN 7.35
[13] ET-Network(No Attention) No 7.79
Ours CRNN No 6.66 ± 0.18
[14] ViT+T5+GPT2 No 17.0
[21] CNN-Transformer (pre-trained) Yes 6.2
[13] ET-Network(with Attention) Yes 5.70
[28] Conv-Transformer No 6.4
Ajami (Fulfulde/Hausa) [40] [40] Pre-trained models from OPEN ITI No 64–84
Ours CRNN No 10.66± 0.04
5.2 Label Error Analysis
Using the CER-HV process described in Section 3.3, we analyzed label and content errors across all six datasets and across all dataset splits (training, validation, and test). The analysis reveals that the presence, type, and prevalence of labeling issues vary substantially depending on both the dataset and the split. In total, five datasets contained mislabeled or corrupted samples. PHTD did not exhibit any detectable labeling errors and is therefore excluded from the subsequent error analysis. For the remaining datasets, entries marked with “—” indicate that the corresponding error type was not observed for that dataset and spli
To evaluate how accurately the proposed CER-based ranking identifies mislabeled samples, we measured the precision of the noise detector on the top 50 highest-ranked samples in each split. While potential errors were flagged across all datasets, precision is reported only for datasets with a sufficiently large number of flagged samples—namely Muharaf, PHTI, and Ajami. The remaining datasets (KHATT and NUST-UHWR) contain fewer than 50 flagged samples per split, making the top-50 precision metric not applicable. Table 3 reports the proportion of flagged samples that were confirmed by human reviewers to contain content errors.
The detector achieves very high precision on Muharaf, with 75% precision in the training split, 70% in validation, and 90% in the test split, reflecting the strong correspondence between high CER and true annotation issues in this historical corpus. PHTI also shows consistently high precision across splits (88% in training, 86% in validation, and 80% in testing), whereas Ajami exhibits lower but still substantial precision (71% in training, 70% in validation, and 68% in testing), which aligns with its visually diverse writing styles and challenging calligraphic forms. Overall, these results demonstrate that CER-based ranking effectively identifies problematic samples in noisy datasets, substantially reducing the amount of manual verification required.
Table 3: Precision of the CER-based noise detection computed over the top 50 highest-ranked samples in each split for datasets with sufficient numbers of flagged samples. Precision reflects the proportion of flagged samples that were confirmed as true label errors; the remaining cases correspond to valid but difficult samples.
Dataset Precision (%)
Train Validation Test
Muharaf 75 70 90
PHTI 88 86 80
Ajami 71 70 68
Table 4 summarizes the distribution of detected error categories across the training, validation, and test splits for all datasets. The results show clear dataset-specific patterns. In KHATT and NUST-UHWR, errors are rare and largely limited to isolated transcription mistakes. In PHTI, transcription and orientation errors dominate across splits. The Muharaf dataset contains a higher proportion of segmentation errors and script-mismatch cases, particularly in the training data, while Ajami exhibits a mixture of transcription, segmentation, orientation, and script-mismatch errors across all splits. These differences reflect variations in dataset construction processes, source material, and annotation pipelines, underscoring the need for dataset-specific quality assessment.
Table 4: Distribution of detected label errors by type across training, validation, and test splits based on the CER-HV annotation process.
Dataset Split
Transcription
Error
Segmentation
Error
Orientation
Error
Script-
Mismatch
Irrelevant
Content
Total
KHATT Train 40 4 44 (0.9%)
Val 5 5 (0.05%)
Test (—)
Muharaf Train 42 99 4 998 288 1431 (6.4%)
Val 20 17 54 10 101 (9.4%)
Test 9 23 7 70 17 126 (9.4%)
PHTI Train 220 2 98 320 (1.3%)
Val 59 2 15 76 (1.4%)
Test 48 4 20 72 (1.3%)
NUST-UHWR Train 3 3 (0.03%)
Val 4 4 (0.04%)
Test 5 5 (0.5%)
Ajami Train 68 47 54 196 6 371 (8.5%)
Val 13 10 15 43 2 83 (8.9%)
Test 20 10 10 44 2 86 (9.2%)
5.3 Qualitative Examples
To illustrate the major error categories identified by the proposed framework, we present representative visual examples. Figure 5 shows examples from each error type: transcription error, segmentation error, orientation error, script mismatch, and irrelevant or non-text content, highlighting the diversity of labeling challenges encountered across the Arabic-script datasets. While non-textual elements such as stamps or signatures and text written in non-target scripts are treated as errors in our HTR-focused setting, it is important to note that such elements may be meaningful or even desirable targets in other research contexts, such as writer identification, archival studies, or multilingual document analysis.
Figure 5: Representative examples of flagged samples from the major error categories identified through the CER-HV pipeline: (a) transcription error, (b) segmentation error, (c) orientation error involving a 180° rotation, (d) orientation error involving diagonally oriented text, where a red arrow indicates the intended reading direction, (e) script mismatch, and (f) irrelevant or non-text content. For transcription errors (a), the model prediction is shown to illustrate the mismatch with the ground truth. For structural errors where correction is straightforward (b–c), the corrected image is displayed. Dataset name, image ID, CER, and ground-truth text are shown for all samples. Examples are drawn from multiple datasets (Muharaf, NUST-UHWR, PHTI, and Ajami), highlighting the diverse manifestations of labeling challenges across Arabic-script handwriting corpora.
In addition to noisy samples, the noise detector identified a subset of correctly labeled but visually challenging examples. These “valid but hard” cases often involve text containing numbers or special symbols (e.g., slashes), which are underrepresented in the training data, as well as samples with low visual clarity, unusually long text lines, heavy stroke overlap, or dense diacritics. Although these samples were not corrected or removed, they provide insight into where the model still struggles and help distinguish genuine annotation errors from intrinsic recognition difficulty. Figure 6 illustrates representative cases.
Figure 6: Representative examples of “valid but hard” samples flagged by the detector. These cases are correctly labeled but visually challenging.
5.4 Quantitative Impact of Label Errors
Having identified the main categories of label errors, we now examine how correcting these errors affects model performance. Using the proposed CER-HV framework, we detected mislabeled or corrupted samples across all data splits and evaluated the impact of excluding these samples. We then analyzed how training on cleaned data further improves recognition performance. The combined effects of evaluation-set and training-set cleaning are summarized in Table 5.
Table 5: Impact of training and evaluation set refinement on CER. The $\Delta$ column represents the cumulative improvement compared to the original noisy baseline (first row) of each dataset.
Dataset Training Set Evaluation Set Validation Split Test Split
(Val/Test) CER (%) $\Delta$ CER (%) $\Delta$
KHATT Raw (Noisy) Raw (Noisy) 8.53 8.45
Raw (Noisy) Cleaned 8.14 -0.39
Cleaned Cleaned 7.93 -0.60 8.3 -0.10
Muharaf Raw (Noisy) Raw (Noisy) 9.82 10.11
Raw (Noisy) Cleaned 8.78 -1.04 8.45 -1.66
Cleaned Cleaned 8.61 -1.21 8.25 -1.86
PHTI Raw (Noisy) Raw (Noisy) 8.37 8.26
Raw (Noisy) Cleaned 7.39 -0.98 7.46 -0.80
Cleaned Cleaned 7.29 -1.08 7.36 -0.90
NUST-UHWR Raw (Noisy) Raw (Noisy) 6.89 6.66
Raw (Noisy) Cleaned 6.57 -0.32 6.43 -0.23
Cleaned Cleaned 6.57 -0.32 6.43 -0.23
Ajami Raw (Noisy) Raw (Noisy) 10.32 10.66
Raw (Noisy) Cleaned 9.50 -0.82 9.42 -1.24
Cleaned Cleaned 9.01 -1.31 8.82 -1.84
Cleaning the evaluation splits results in noticeably reduced CER across all datasets. This improvement is relatively small (0.2–0.5% absolute) for the cleanest datasets, such as KHATT and NUST-UHWR, which exhibit minimal annotation noise. Moderate gains of around 1% are observed for PHTI, while the largest improvements, in the range of 1–1.6%, are recorded for Muharaf and Ajami. These cleaned evaluation splits therefore serve as revised benchmarks for the corresponding datasets, and we release them publicly to support more reliable and reproducible evaluation in future HTR research. To assess how label noise affects the learned model representations, we applied the CER-HV procedure to the training splits and retrained the CRNN on the resulting cleaned datasets $D^{\prime}$. By comparing the second and third rows for each dataset in Table 5, we can isolate the gain attributable specifically to training on cleaned data, beyond the effect of evaluation-set refinement.
For KHATT and NUST-UHWR, the proportion of noisy training samples is below 1%, consisting primarily of isolated transcription errors. In these settings, retraining on cleaned data yields negligible additional improvement. This outcome is expected, as such low noise levels are unlikely to bias the learned representation and may even act as a mild regularizing factor during training.
A similar pattern is observed for Muharaf. Although the overall fraction of flagged training samples is higher, a substantial portion corresponds to script-mismatch cases. While these samples are out of scope for Arabic recognition, they do not necessarily induce incorrect character-level supervision in the same way as segmentation or orientation errors, and their removal therefore has limited impact on the learned model. The small CER differences observed after retraining (on the order of $\pm$ 0.2%) fall within the range commonly attributed to stochastic variation in neural network optimization rather than a fundamental change in model capacity.
In contrast, the Ajami dataset provides clearer evidence of the benefit of training-side cleaning. With the highest noise density in the training split (8.5%), Ajami exhibits a noticeable additional gain when retrained on cleaned data, with validation CER improving from 9.50% to 9.01%. Unlike the other datasets, Ajami contains a higher proportion of structurally disruptive noise sources, such as segmentation and orientation errors, which directly affect the visual–text alignment learned by the model. Removing these samples during training therefore leads to a more consistent supervision signal and improved recognition performance.
Taken together, these results show that even limited amounts of label noise can substantially distort model assessment, emphasizing the need for systematic validation of evaluation sets to ensure reliable and fair benchmarking in HTR. While evaluation-set cleaning is critical for obtaining trustworthy performance estimates, training-set cleaning provides measurable benefits primarily when noise levels are sufficiently high and dominated by structurally disruptive errors that interfere with sequence learning. In settings where training noise is sparse or mild, retraining on cleaned data yields diminishing returns.
6 Conclusion and Future Work
This work shows that data quality is a major factor in HTR for Arabic-script languages, yet it has received far less attention than model design. By combining a carefully configured CRNN with the CER-based scoring used in the CER-HV framework, we revealed a wide range of label and content issues across multiple Arabic-script datasets. Cleaning these errors improved evaluation results in every case, confirming that part of the difficulty reported in Arabic-script HTR stems from inconsistent or noisy labels rather than from script complexity alone. Importantly, CER-HV represents a task-specific reformulation of learning-dynamics-based noise detection that makes such approaches viable for CTC-based text recognition.
The results also demonstrate that a well-configured CRNN can achieve state-of-the-art performance on several datasets without synthetic data or Transformer attention. This highlights the importance of strong baselines and shows that deeper or more complex models are not always required to reach competitive accuracy. For example, on KHATT the CRNN achieves a CER of 8.45%, matching or surpassing recent, more complex systems, while on the historical Muharaf dataset it improves substantially over the published baseline when evaluated on the public split.
The effectiveness of the human-in-the-loop component in CER-HV is closely tied to the characteristics of handwritten text recognition. In CTC-based models, loss values are an indirect and often unreliable indicator of label correctness, whereas CER directly reflects transcription quality and is immediately interpretable by human reviewers. Because line-level HTR datasets typically contain thousands rather than millions of samples, restricting inspection to a small subset of high-CER samples makes targeted verification feasible and low cost. This lightweight process allows obvious structural errors—such as segmentation failures, orientation mistakes, script mismatch, or non-text content—to be identified efficiently without exhaustive manual review.
Beyond its use in analyzing existing benchmarks, CER-HV can serve as a practical dataset validation tool. Applied during dataset creation or prior to benchmarking, it enables researchers to identify likely mislabeled or out-of-scope samples and verify only a small number of high-risk cases. Our findings further highlight that data quality represents a complementary dimension to architectural innovation: when annotation issues dominate, increasing model complexity does not reliably improve performance and may instead reinforce incorrect supervision.
A limitation of the present study suggests directions for future work. First, cleaning was implemented through exclusion of noisy samples rather than correction. While this strategy enables a controlled analysis of how label noise affects measured recognition accuracy, it does not exploit the full information contained in the identified samples. Different error types would require different remedies, such as pseudo-labeling for transcription errors, dedicated segmentation models for segmentation errors, or automatic rotation strategies for orientation errors. The released CER-HV error annotations and trained recognition models provide a practical foundation for active learning and semi-automated correction methods in future research.
Although this study focuses on HTR, the underlying two-stage procedure is applicable to other data-limited recognition problems where annotation is costly and errors strongly affect evaluation outcomes. In domains such as medical imaging or symbol detection in engineering drawings, datasets are often small enough that targeted human verification is feasible. Extending CER-HV to such settings, together with broader studies across different architectures, including Transformer-based recognizers, represents an important avenue for future work.
Overall, these findings show that progress in Arabic-script HTR depends not only on advanced models but also on reliable data, and that systematic dataset quality assessment provides a straightforward path toward stronger baselines, fairer evaluation, and more reproducible research.
7 Acknowledgements
This research is financially supported by the European Regional Development Fund and the MARTINA-project (no. 20367152).